OpenStudy (anonymous):

16. Jennifer writes each of the letters from A to Z on a separate card and places all the cards in a box. She randomly draws three cards from the box. What is the probability that all three cards are vowels?

OpenStudy (anonymous):

prob of first = vowel = 5/26 2nd = 4/25 3rd = 3/24 reqd probability = 5/26 * 4/25 * 3/24

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

yw